An automobile radiator contains 16L of antifreeze and water. This mixture is 30% antifreeze. How much of this mixture should be drained and replaced with pure antifreeze so that there will be 50% antifreeze? (Not from a textbook.) I'm ok to solve is just adding antifreeze, but can't figure out how to set it up so that the solution is drained and replaced with pure antifreeze.
Answer by scott8148(6628)

you're going to drain xL of 30% antifreeze leaving (16-x)L ... to this you'll add xL of 100% antifreeze to get 16L of 50%

.3(16-x)+x=.5(16) ... 4.8-.3x+x=8 ... .7x=3.2 ... x=32/7 or about 4 1/2 L